# Recurrent Inference Machine for Medical Image Registration

### v1

@@ 4.1 Experiment 1: Comparison Study with Varying Data Availability
The results for OASIS are depicted in Fig 4(a), as well as an illustrative visualization of RIIR inference on an example test data, as shown in Fig. 5. It is evident that RIIR outperforms when data availability is severely limited and maintains consistent performance across various data availability scenarios, showcasing its data efficiency and accuracy. For a more detailed comparison, the statistics when full data availability is guaranteed are presented in Tab. 1 for OASIS.
(a) Dice score on OASIS.
(b) $\mathcal{D}_{\text{PCA1}}$ on mSASHA.
Figure 4: Results of Experiment 1 with boxplots for Dice score and $\mathcal{D}_{\text{PCA1}}$ on two datasets, respectively. The circle denotes the mean of the metric of interest. The segmentation metric Dice is calculated for all 35 segmentation labels and post-processed by taking the average. The group-wise metric $\mathcal{D}_{\text{PCA1}}$ was calculated based on further center-cropping at a ratio of $70\%$ on the warped images.
Table 1: Results of Experiment 1 on OASIS with $100\%$ data availability. The metrics are reported with a format of mean(standard deviation). A double-sided Wilcoxon test is calculated for Dice score, and denotes a significant difference compared with the baseline VoxelMorph in terms of Dice score ($p<0.05$).
Model Dice $\uparrow$ HD $\downarrow$ $\text{std}(\log|J|)$downarrow
VoxelMorph 0.694(0.035) 3.78(0.54) 0.26(0.11)
GraDIRN 0.697(0.034) 3.76(0.54) 0.31(0.14)
RIIR 0.724(0.031) 3.71(0.53) 0.32(0.12)
The results of this experiment on mSASHA are shown in Fig. 4(b) using a composition of boxplots. An illustrative evolving visualization of RIIR inference demonstrated in a step-wise fashion can be found in Fig. 6. The corresponding boxplots for $\mathcal{D}_{\text{PCA2}}$ can be found in the supplementary material. Similarly, the statistics of mSASHA when full data availability is guaranteed are shown in in Tab. 2.
Table 2: Results of Experiment 1 on mSASHA with $100\%$ data availability. The metrics are reported with a format of mean(standard deviation). A double-sided Wilcoxon test is calculated for $\mathcal{D}_{\text{PCA1}}$, and denotes a significant difference compared with the baseline VoxelMorph ($p<0.05$).
Model $\mathcal{D}_{\text{PCA1}}$downarrow $\mathcal{D}_{\text{PCA2}}$downarrow $\text{std}(\log|J|)$downarrow
VoxelMorph 0.425(0.126) 35.88(20.13) 0.33(0.25)
GraDIRN 0.343(0.078) 34.89(12.69) 1.56(0.78)
RIIR 0.324(0.072) 34.49(11.99) 2.18(0.79)
Overall, both iterative methods performed better across most data availability scenarios on mSASHA. As data availability approached saturation, the performances of both iterative methods converged. A further qualitative comparison among the three models is shown in Fig. 7.
Figure 7: Qualitative comparison of mSASHA. The images were scaled with the same color range in each row for comparison. Artifacts (highlighted by red arrow) in the myocardium region are observed for both VoxelMorph and GraDIRN.
@@ 4.2 Experiment 2: Inclusion of Hidden States
The architectural settings were kept the same as in the aforementioned experiments and results are shown in Fig. 12. In addition to better performance in the two datasets, we also empirically noticed that the training was more stable with two hidden states both activated.
Figure 8: Results of Experiment 2 evaluated on OASIS validation set regarding Dice score. Two-sided Wilcoxon tests were conducted for $[1,1]$ against other settings with statistical significance $(p<0.05)$, except for $[0,1]$ $(p=0.11)$.
Figure 9: Results of Experiment 2 evaluated on mSASHA validation set regarding $\mathcal{D}_{\text{PCA1}}$. Here, for example, $[0,0]$ denotes the case that no hidden states are considered, and $[1,1]$ denotes both hidden states were considered in the pipeline. Two-sided Wilcoxon tests were conducted for $[1,1]$ against other settings with statistical significance $(p<0.05)$.
4.3 Experiment 3: Choice of Inner Loss Functions $\mathcal{L}_{\text{inner}}$
Figure 10: Results of Experiment 3 evaluated on mSASHA validation set. Raw denotes no registration is implemented to the image series.
The quantitative results, measured by $\mathcal{D}_{\text{PCA1}}$, are shown in Fig. 10. It is noteworthy that the performance when considering NMI as the likelihood function did not align with the performance of the other two scenarios. In response, we conducted a further investigation, presenting a qualitative visualization of $\nabla\mathcal{L}_{\text{inner}}$ in Fig. 11 to shed light on this discrepancy. In our implementation, autograd (Paszke et al., 2017) were used to calculate $\nabla\mathcal{L}_{\text{inner}}$ for NCC and NMI, while we used the analytical gradient of MSE due to its simplicity. However, the Gaussian approximation of NMI involves flattening $I_{\text{fix}}$ and $I_{\text{mov}}\circ\bm{\phi}$ to compute the joint probability, potentially resulting in the scattering pattern shown in Fig. 11 and influencing performance, as $\nabla\mathcal{L}_{\text{inner}}$ is the sole image information passed into RIIR cell.
Figure 11: Qualitative check of Experiment 3 with an example pair from mSASHA. Three gradient images are taken from the last step of the RIIR inference pipeline at the same epoch.
4.4 Experiment 4: Inclusion of Inner Loss Gradient as RIIR Input
The results are shown in Fig. 9. It can be observed that the network struggled if the warped image $I_{\text{mov}}\circ\bm{\phi}$ is only implicitly fed to the RIIR cell, i.e., there’s an additional cost to learn the transformation. During our training, we also noticed that when trained with explicit warped images, the network tends to have a larger deformation magnitude since the images are more detailed compared to the gradient input (as shown in Fig. 11). Based on the results, we conjecture that the explicit input often offers more details to guide the optimization which benefits the training. However, it can be also redundant for training when data availability is limited thus potentially leading to overfitting.
Figure 12: Results of Experiment 4 evaluated on mSASHA validation set. In the experiment, we have Implicit: $[\bm{\phi}_{t},I_{\text{mov}},I_{\text{fix}}]$; Explicit: $[\bm{\phi}_{t},I_{\text{mov}}\circ\bm{\phi}_{t},I_{\text{fix}}]$; and Default: $[\bm{\phi}_{t},\nabla_{\bm{\phi}_{t}}\mathcal{L}_{\text{inner}}]$.
@@ 4.5 Experiment 5: RIIR Architecture Ablation
We here demonstrate the model architecture ablation by showing the corresponding boxplots in Fig. 13. Apart from the main experiments shown previously, these experiments can be regarded as minor ablation studies, aiming to strike a balance between computational precision and inference speed.
Figure 13: Results of Experiment 5. For the ablation study on network steps, double-sided Wilcoxon tests suggest no significant difference is found for $t=6,8,10$. While for the ablation on loss weighting strategy, the exponential decay (denoted by ‘Decay’) shows a significant difference to uniform weighting (‘Uniform’) $(p<0.05)$.
@@ 5 Discussion
In this study, we introduced RIIR, a deep learning-based medical image registration method that leverages recurrent inferences as a meta-learning strategy. We extended Recurrent Inference Machines (RIMs) to the image registration problem, which has no explicit forward models. RIIR was extensively evaluated on public brain MR and in-house quantitative cardiac MR datasets, and demonstrated consistently improved performance over established deep learning models, in both one-step and iterative settings. Additionally, our ablation study confirmed the importance of incorporating hidden states within the RIM-based framework.
The acclaimed performance improvement of registration is especially pronounced in scenarios with limited training data as demonstrated in Fig. 4. Notably, RIIR achieved superior average evaluation metrics on segmentation accuracy and groupwise alignment, with significantly lower variance. This performance advantage was pronounced in the mSASHA experiments with very limited training data, a stark contrast to the baseline models which required much more training data to reach similar performances. Though both GraDIRN and RIIR are iterative methods, GraDIRN isolates the update of explicit $\mathcal{L}_{\text{sim}}$ and deep learning-based $\mathcal{L}_{\text{reg}}$ with no internal states, thus potentially resulting in worse generalization and slower convergence compared with RIIR. Notably, GraDIRN initialized the deformation field randomly by default, which could lead to optimization difficulties when data were extremely limited during training. Furthermore, in the mSASHA dataset, the qualitative visual comparison in Fig. 7 highlights that incorporating gradient information in our RIIR method can better preserve the original structure of the myocardium in the registration process. Furthermore, the step-wise loss function, as used in Eq. 13, also prevents exaggerated registration that can compromise the accuracy and affect the resulting quantitative maps as shown in the baseline methods.
@@
Although hidden states were used in the original RIM and later works (Putzky and Welling, 2017; Putzky et al., 2019; Sabidussi et al., 2021, 2023), it has not been investigated in detail regarding its impact on optimization of RIM-based methods. Our second experiment investigates the impact of these hidden states within RIIR. Our findings reveal that the presence of hidden states, as proposed in the original RIM work (Putzky and Welling, 2017), contributes positively to the performance of our model, as shown by the quantitative results in Fig. 9 and Fig. 8. However, in the 3D OASIS dataset, this improvement was less pronounced compared to the mSASHA dataset, where multiple images in a time series are evaluated together.
We further investigated the influence of inner loss selection (Experiment 3), which was not studied before in iterative deep learning-based registration (Qiu et al., 2022). In mSAHSA, with $\mathcal{L}_{\text{sim}}$ chosen as NMI in $\mathcal{L}_{\text{outer}}$, the quantitative results are not as satisfactory as other choices (MSE, NCC). This is probably due to the implementation of NMI operation that may dissect the neighborhood structure of the original image for RIIR Cell as input as shown in Fig. 11. Nevertheless, this discrepancy suggests that the gradient information of $\mathcal{L}_{\text{inner}}$ can be considered as a form of compressed image data that is processed with the convolutional nature in RIIR Cell. The ablation study on input combinations (Experiment 4) indicates that, in scenarios with limited data, the inclusion of gradient input effectively prevents model overfitting, thus leading to better generalization performance as shown in Fig. 12.
RIIR’s superior registration performance and data efficiency suggest its potential for applications in medical image registration. However, it is necessary to acknowledge the current limitations, to further enhance the framework in future work. First, the performance of RIIR could be bottlenecked by the simplistic convolutional GRU unit, where potential improvement can be considered such as dilated convolutional kernels which can preserve larger receptive fields. Second, the expanding computation graph of RIM-based models, with increasing inference steps, leads to heavy GPU memory consumption.
@@ Acknowledgments
This work was partly supported by the TU Delft AI Initiative, Amazon Research Awards, and the National Heart, Lung, and Blood Institute, National Institutes of Health by the Division of Intramural Research (Z1A-HL006214).
Ethical Standards
The work follows appropriate ethical standards in conducting research and writing the manuscript, following all applicable laws and regulations regarding treatment of animals or human subjects.
Conflicts of Interest
We declare we don’t have conflicts of interest.
A Supplementary Material
(a) HD on test OASIS data.
(b) $\mathcal{D}_{\text{PCA2}}$ on test mSASHA data.
Figure 14: Supplementary results of Experiment 1.
We show the boxplots of Experiment 1 here in terms of HD and $\mathcal{D}_{\text{PCA2}}$ on OASIS and mSASHA datasets respectively in Fig. 14.

### v2

@@ 4.1 Experiment 1: Comparison Study with Varying Data Availability
An illustrative visualization of RIIR inference on an example test data, can be found in Fig. 5. The results for OASIS are presented in Fig 6, as well as . LapIRN, leveraging its multi-resolution architecture, also demonstrates robust performance. It is evident that RIIR outperforms most deep learning-based baselines when data availability is severely limited and maintains consistent performance across various data availability scenarios, showcasing its data efficiency and accuracy.
Figure 6: Results of Experiment 1 with boxplots for Dice score on OASIS. The circle denotes the mean of the metric of interest. The segmentation metric Dice is calculated for all 35 segmentation labels and post-processed by taking the average.
Figure 7: Results of Experiment 1 with boxplots for TRE on NLST. The circle denotes the mean of the metric of interest. The TRE is calculated based on the anatomically meaningful keypoint pairs provided by corrField algorithm within lung lobes. For visualization, outliers over a TRE of 20mm were excluded but used for statistical calculation.
Figure 8: Results of Experiment 1 with boxplots for $D_{pca1}$ on mSASHA. The circle denotes the mean of the metric of interest. The group-wise metric $D_{\text{PCA1}}$ was calculated based on further center-cropping at a ratio of $70\%$ on the warped images.
The results for NLST are shown in Fig. 7. An illustrative visualization of RIIR inference on an inhale-exhale lung CT pair, can be found in Fig. 20. The registration of lung CT presents unique challenges due to the large deformation between respiratory phases. RIIR demonstrates superior performance in capturing these large deformations, achieving lower TRE while maintaining anatomically plausible transformations.
Figure 9: Results of Experiment 2 evaluated on OASIS validation set (left) regarding Dice score and NLST validation set (right) regarding TRE. Here, for example, $[0,0]$ denotes the case that no hidden states are considered, and $[1,1]$ denotes both hidden states were considered in the pipeline. Two-sided Wilcoxon tests were conducted for $[0,1]$ against other settings with statistical significance $(p<0.05)$, except for $[0,0]$ $(p=0.076)$ in OASIS dataset.
The results of this experiment on mSASHA are shown in Fig. 8 using a composition of boxplots. Both LapIRN and RCVM achieved superior performance in group-wise registration, with RCVM showing slightly better results in terms of $\mathcal{D}_{\text{PCA1}}$. Our proposed RIIR demonstrated comparable performance levels in terms of $\mathcal{D}_{\text{PCA1}}$. The qualitative visualization of RIIR inference on mSASHA test split is shown in Fig. 21.
For a comprehensive quantitative comparison, Table 1 presents detailed statistics across all datasets under full data availability, including performance metrics, model parameters, memory consumption, and computational time.
@@ 4.2 Experiment 2: Inclusion of Hidden States
The architectural settings were kept the same as in the aforementioned experiments, and the results are shown in Fig. 9. Although improved performance using hidden states ([0, 1]) over no hidden states is only observed in the NLST dataset, we empirically noticed that training was more stable when hidden states were enabled. It is also notable that the inclusion of hidden states in the pipeline does not incur significant computational overhead. For the OASIS dataset, the addition of hidden states only increases VRAM consumption by approximately 800MB, while maintaining the model’s computational speed.
Table 2: Comparison of Explicit Input and RIIR under 5% data availability in Experiment 3. For each metric, we report mean±$\pm$±std.
Dataset $|J_{\phi}|_{\leq 0}(\%)$ std$(\log|J_{\phi}|)$ $\operatorname{mag}(\nabla|J_{\phi}|)$
OASIS
Explicit $0.014\pm 0.015$ $0.299\pm 0.133$ $0.035\pm 0.003$
RIIR $0.019\pm 0.012$ $0.350\pm 0.088$ $0.034\pm 0.003$
NLST
Explicit $0.092\pm 0.108$ $0.575\pm 0.372$ $0.057\pm 0.003$
RIIR $0.085\pm 0.084$ $0.581\pm 0.309$ $0.055\pm 0.004$
4.3 Experiment 3: Inclusion of Inner Loss Gradient as RIIR Input
The results are shown in Fig. 10. It can be observed that the network struggled if the warped image $I_{\text{mov}}\circ\bm{\phi}$ is only implicitly fed to the RIIR cell, i.e., there’s an additional cost to learn the transformation. When training data is abundant (100% availability), RIIR and Explicit input achieve comparable performance, as evidenced by the non-significant statistical differences ($p=0.17$ for OASIS and $p=0.27$ for NLST). This suggests that with sufficient training data, both approaches can effectively learn the registration task, though RIIR maintains its advantage over other input types. However, RIIR input demonstrates more robust performance in data-limited scenarios, particularly when only 5% of the training data is available. We also found that for both datasets when data availability is $5\%$, the deformation-related metrics have smaller variation for RIIR input compared with Explicit inputs as shown in Table 2.
Figure 10: Results of Experiment 3 evaluated on OASIS validation set (left) and NLST validation set (right). For OASIS with 5% data availability, RIIR input shows significance over all types except RIM ($p=0.81$). At 100%, significance remains except for Explicit input ($p=0.17$). For NLST with 5% data availability, RIIR input shows significance over all types. At 100%, significance remains except for Explicit input ($p=0.27$).
4.4 Experiment 4: Regularization Analysis
The results of regularization analysis on OASIS and NLST datasets are shown in Fig. 11 and Fig. 12, respectively. For OASIS, diffusion regularization achieves better Dice scores compared to affine registration across all weights. For NLST, while all three regularization terms improve the TRE over affine registration, curvature regularization exhibits a higher percentage of negative Jacobian determinants. For elastic regularization, it is worth noting that different parameter settings for $\lambda_{e}$ and $\mu$ were used across datasets to reflect tissue differences, which may lead to varying regularization effects.
Figure 11: Results of regularization analysis on OASIS validation set. Different regularization weights are compared in terms of Dice score. Here w/o reg denotes the results without regularization and negJ denotes the percentage of negative Jacobian determinant.
Figure 12: Results of regularization analysis on NLST validation set. Different regularization weights are compared in terms of TRE. Here w/o reg denotes the results without regularization and negJ denotes the percentage of negative Jacobian determinant.
@@ 4.5 Experiment 5: RIIR Architecture Ablation
Here we demonstrate the model architecture ablation by showing the corresponding boxplots in Fig. 13. The increasing number of steps leads to proportionally higher VRAM consumption and inference time. Specifically, VRAM usage ranges from approximately 8GB (4 steps) to 24GB (12 steps), with corresponding inference times varying from 0.40s to 1.12s. Apart from the main experiments shown previously, this experiment can be regarded as a minor ablation study, aiming to strike a balance between computational precision and inference speed.
Figure 13: Results of Experiment 5. For the ablation study on network steps on OASIS, two-sided Wilcoxon tests suggest significant difference is found for $t=6$ against all other scenarios.
@@ 5 Discussion
In this study, we introduced RIIR, a deep learning-based medical image registration method that leverages recurrent inferences as a meta-learning strategy. We extended Recurrent Inference Machines (RIMs) to the image registration problem, which has no explicit forward models. Given the absence of explicit forward models, our approach can be viewed as a case of amortized iterative inference, where the network learns to progressively refine the registration. RIIR was extensively evaluated on public brain MR, lung CT, and in-house quantitative cardiac MR datasets, and demonstrated consistently improved performance over established deep learning models, both in one-step and iterative settings. Additionally, our ablation study confirmed the importance of incorporating hidden states within the RIM-based framework.
The acclaimed improvement in registration performance is especially pronounced in scenarios with limited training data, as demonstrated in Figs. 6, 7, and 8. In particular, RIIR achieved superior average evaluation metrics with lower variance. This performance advantage was particularly pronounced in the NLST experiments, where large deformations need to be estimated with very limited training data. Most baseline models required substantially more training data to achieve comparable performance levels, with the exception of L2O, which benefits from additional input modalities such as dynamically sampled coordinates and MIND features. Though both GraDIRN and RIIR are iterative methods that use similarity gradients, GraDIRN isolates the update of explicit $\mathcal{L}_{\text{sim}}$ and deep learning-based $\mathcal{L}_{\text{reg}}$ with no internal states, potentially resulting in worse generalization and slower convergence compared to RIIR. In particular, GraDIRN initialized the deformation field randomly by default, which could lead to optimization difficulties when the data were extremely limited during training. RCVM achieves lower folding rates ($|J_{\phi}|_{\leq 0}$) through its composed deformation strategy, where the final transformation is obtained by composing multiple smaller deformations. Although this approach effectively prevents topology-breaking deformations, it comes at the cost of an increased number of parameters. While the training and inference times on the OASIS dataset are the second longest among learning-based methods (after L2O), RCVM exhibits comparable runtime with RIIR for both 3D datasets, and shows faster inference time on the 2D dataset (mSASHA). Notably, LapIRN, as a single-pass multi-resolution method, shows remarkable advantages in both VRAM efficiency and inference speed.
@@
Although hidden states were used in the original RIM and later work (Putzky and Welling, 2017; Putzky et al., 2019; Sabidussi et al., 2021, 2023), their impact on the optimization of RIM-based methods has not been investigated in detail. Our second experiment investigates the impact of these hidden states within RIIR. Our findings reveal that the presence of hidden states, as proposed in the original RIM work (Putzky and Welling, 2017), contributes positively to the performance of our model, as shown by the quantitative results in Fig. 9. Unlike L2O which operates on full resolution features at the output, our implementation of hidden states at downsampled resolution leads to minimal additional memory overhead.
The ablation study in input combinations (Experiment 3) demonstrates that, in scenarios with limited data, RIIR with images gradient input achieves superior registration performance in anatomical evaluation metrics, as shown in Fig. 10. This improvement is particularly evident in the NLST dataset, where RIIR achieves lower TRE and smoother deformation fields. Although including gradient input can be considered to offer additional information, its impact on regularization varies between datasets (Table 2). This could possibly be due to the different similarity losses and registration objectives (intra-subject for NLST versus inter-subject for OASIS).
The regularization analysis reveals that RIIR’s performance remains dependent on both the choice and weight of regularization terms. As shown in Fig. 11 and Fig. 12, different regularization strategies lead to varying trade-offs between registration accuracy and deformation regularity. This suggests potential future improvements by incorporating adaptive regularization schemes (Hoopes et al., 2021; Mok and Chung, 2021; Reithmeir et al., 2024). Such extensions could enhance RIIR’s robustness across different clinical scenarios without manual parameter tuning.
The superior registration performance and data efficiency of RIIR suggest its potential for applications in medical image registration. However, it is necessary to acknowledge the current limitations, to further enhance the framework in future work. From an architectural perspective, RIIR employs a relatively conventional design that lacks multi-resolution capability, which has proven effective in methods like LapIRN. The simple convolutional GRU structure could also be enhanced with modern components such as dilated convolutions for larger receptive fields or attention mechanisms for better feature extraction. Another significant limitation lies in GPU memory consumption. Despite having fewer parameters than LapIRN and RCVM, RIIR requires more VRAM due to its recurrent nature, and this memory usage increases linearly with the number of inference steps, as demonstrated in Experiment 5. To better adapt to downstream tasks, potential improvements could include semi-supervised strategy, instance optimization and adaptive regularization strategies, which could enhance the flexibility of the model in different clinical scenarios.
@@ Acknowledgments
This work was partly supported by the TU Delft AI Initiative, Amazon Research Awards, the Dutch Research Council (NWO), and the National Heart, Lung, and Blood Institute, National Institutes of Health by the Division of Intramural Research (Z1A-HL006214).
Supplementary Materials
Additional Results of Experiment 1
We show the boxplots of Experiment 1 here in terms of HD on OASIS dataset in Fig. 14. Additionally, we present the Dice scores for lung lobes on the NLST dataset in Fig. 15. For mSASHA dataset, we show the $\mathcal{D}_{\text{PCA2}}$ metric in Fig. 16. A qualitative comparison on OASIS is shown in Fig. 19.
Figure 14: Supplementary results of Experiment 1 on test OASIS data, showing Hausdorff Distance (HD) comparison between different methods. Lower HD values indicate better registration accuracy and structural preservation.
Figure 15: Supplementary results of Experiment 1 on test NLST data, showing Dice scores for lung lobes between different methods. Higher Dice scores indicate better registration accuracy in aligning lung structures.
Figure 16: Supplementary results of Experiment 1 on test mSASHA data, showing $\mathcal{D}_{\text{PCA2}}$ metric comparison between different methods.
Implementation study: Choice of Inner Loss Functions $\mathcal{L}_{\text{inner}}$ for mSASHA
Figure 17: Qualitative check of gradient information with an example pair from mSASHA. Three gradient images are taken from the last step of the RIIR inference pipeline at the same epoch.
Figure 18: Results evaluated on mSASHA validation set. Raw denotes no registration is implemented to the image series.
Since it has been proposed previously that NMI is a more suitable choice of $\mathcal{L}_{\text{sim}}$ for $\mathcal{L}_{\text{outer}}$ in qMRI registration (de Vos et al., 2020) over NCC and MSE, we focused on the choice of $\mathcal{L}_{\text{inner}}$ to investigate the registration performance and the information propagated through the RIIR on mSASHA. In our implementation, autograd (Paszke et al., 2017) were used to calculate $\nabla\mathcal{L}_{\text{inner}}$ for NCC and NMI, while we used the analytical gradient of MSE due to its simplicity. However, the NMI calculation requires estimating the joint probability distribution using Parzen window (Gaussian kernel) approximation, which involves flattening $I_{\text{fix}}$ and $I_{\text{mov}}\circ\bm{\phi}$ to compute the joint probability. This approximation results in the scattering pattern shown in Fig. 17 and influence the performance when using the RIM input ($[\bm{\phi}_{t},\nabla_{\bm{\phi}_{t}}\mathcal{L}_{\text{inner}}]$) as shown in Fig. 18.
Figure 19: Qualitative comparison of different registration methods on a representative OASIS test case. The bottom row shows the corresponding deformation fields where ‘Affine’ denotes no further registration applied.
Figure 20: A visualization of RIIR inference on NLST test split, visualized with a 2D slice and in-plane deformation. The inference step was set to $6$ in both training and inference. All images in the same row were plotted using the same color range for better consistency.